A die is rolled five times.Determine the probability of obtaining 3 sixes
Answer by stanbon(75887)

A die is rolled five times.Determine the probability of obtaining 3 sixes.
---
Binomial Problem with n=5 and p = 1/6
-----
P(x=3) = 5C3(1/6)^3*(5/6)^2 = 10*(1/216)(25/36) = 0.03215
------------------------
Cheers,
Stan H.